Find the p-value for the z-test. (Round your answer to four decimal places.)

a right-tailed test with observed z = 1.18

p-value =
Answer #1

solution

z= 1.18

it is  right-tailed test

P(z >1.18 ) = 1 - P(z <1.18 ) = 1-0.881=0.1190

P-value = 0.1190 ( using z table)